OpenStudy (deadshot):

What is the simplified form of x plus 1 over x squared plus x minus 6 divided by x squared plus 5x plus 4 over x minus 2 ?

OpenStudy (anonymous):

can you write it down?

OpenStudy (deadshot):

\[\frac{ (x+1) }{ x^2+x-6 }\div \frac{ x^2+5x+4 }{ (x-2) }\]

OpenStudy (anonymous):

x^2 + 5x + 4 = (x+1)(x+4)

OpenStudy (anonymous):

x^2 + x - 6 = (x-2)(x+3)

OpenStudy (deadshot):

So the answer is \[\frac{ 1 }{ (x+3)(x+4) }\] right?